OpenStudy (anonymous):

Hopefully this is the last one: I can't seem to rationalize this limit to yield a non division by zero scenario. Can anyone help?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 8}\frac{ x-8 }{ \sqrt{x+1}-3}\]

ganeshie8 (ganeshie8):

why not ?

OpenStudy (anonymous):

Well regardless of whether I use the numerator or denominator's conjugate, I still end up with a zero as the denominator.

ganeshie8 (ganeshie8):

lets see :) \(\large \lim_{x \rightarrow 8}\frac{ x-8 }{ \sqrt{x+1}-3} \) rationalize denominator, \(\large \lim_{x \rightarrow 8}\frac{ x-8 }{ \sqrt{x+1}-3} \times \frac{ \sqrt{x+1}+3}{\sqrt{x+1}+3} \)

ganeshie8 (ganeshie8):

\(\large \lim_{x \rightarrow 8}\frac{ x-8 }{ \sqrt{x+1}-3} \) rationalize denominator, \(\large \lim_{x \rightarrow 8}\frac{ x-8 }{ \sqrt{x+1}-3} \times \frac{ \sqrt{x+1}+3}{\sqrt{x+1}+3} \) \(\large \lim_{x \rightarrow 8}\frac{ (x-8)(\sqrt{x+1}+3) }{ (\sqrt{x+1})^2-3^2} \)

ganeshie8 (ganeshie8):

\(\large \lim_{x \rightarrow 8}\frac{ x-8 }{ \sqrt{x+1}-3} \) rationalize denominator, \(\large \lim_{x \rightarrow 8}\frac{ x-8 }{ \sqrt{x+1}-3} \times \frac{ \sqrt{x+1}+3}{\sqrt{x+1}+3} \) \(\large \lim_{x \rightarrow 8}\frac{ (x-8)(\sqrt{x+1}+3) }{ (\sqrt{x+1})^2-3^2} \) \(\large \lim_{x \rightarrow 8}\frac{ (x-8)(\sqrt{x+1}+3) }{ x-8} \)

ganeshie8 (ganeshie8):

next step is obvious ! ? :)

OpenStudy (anonymous):

Yeah, I see. Thanks a lot.

ganeshie8 (ganeshie8):

np :)